SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 3 September 2004

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Friday, September 3, 2004

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Today's Orlan EVA-11 by CDR Padalka and FE/SO Fincke, their fourth spacewalk, was yet another solid success for the ISS program.  Spending 5h 21m outside the Russian segment (RS), the two spacewalkers swapped out a thermal control system fluid control valve panel on the FGB, installed new hardware associated with Europe's ATV (Automated Transfer Vehicle), added holding fixtures, protective covers and "bookend"-type brackets to handrails (to keep Russian tethers from winding themselves around the handholds) and took photographs of external experiment packages.  All objectives were accomplished (in 28 min less time than expected).   [After DC1 airlock hatch #1 opening at 12:43pm EDT, 7 min ahead of plan, Padalka and Fincke egressed at 12:51pm with equipment bundles and translated along handrails to the FGB panel worksite were they removed the old RRZh-1 flow regulator valve panel (1:29pm) and replaced it with a spare unit (1:40pm).  A ground-controlled test later showed that it is working nominally.  They also attached "fairlead" safety tether guides on FGB handrails, held still for a 17-min. "Motionless" test during the first orbital night period to support the ongoing extraneous torque investigation, then returned to the DC1 for temporary stowage of the old 70 kg RRZh (later thrown overboard in retrograde direction prior to ingress) and picking up the next batch of equipment.  Continuing on to the SM aft end after RS thrusters were inhibited and CMG control enabled, they installed three antennas (WAL-1, -2 & -3) for ATV proximity ops, and removed the covers from all five antennas (three more ATV prox ops antennas will be installed next year by Chiao and Sharipov of Expedition 10).  After taking more photographs, they returned to the DC-1 for more handrail work, with attitude control being returned to RS thrusters (5:32pm).  The crew ingressed "Pirs" at ~6:00pm.  Hatch closure occurred at 6:04pm, after 5h 21m.  This was the 56th EVA in support of ISS assembly/maintenance, the 31st from the station itself, the sixth for Gennady Padalka and the fourth for Mike Fincke.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Earlier, after crew wake-up at 3:45am (1h 45m later than usual), all pre-EVA activities had proceeded smoothly and on schedule, starting out with CDR Padalka and FE Fincke undergoing another MO-9 urine biochemistry test.   [A second session with the Urolux equipment will be conducted by both crewmembers immediately after the EVA.  The CDR then stows the gear.]

Meanwhile, final ground-commanded station preps for unmanned mode and EVA ops were underway.   [These included contingency telemetry downlink tests, shell heater ops, TCS (thermal control system) Loop B PCVP (pump & control valve package) powerdown, PCU-2 (plasma contactor unit #2) activation (joining the already discharging PCU-1), TCCS (trace contaminant control subsystem) deactivation, configuring S-band, etc.]

Final onboard pre-EVA activities by the crew included:

  • Configuring the Ethernet OpsLAN (Operations Local Area Network) for the RS/USOS hatch closure (and resetting after post-EVA hatch opening);
  • closing U.S. hatches between Lab & Node (2), Node strbd & Airlock, and Node & PMA-1 (at 5:30am), with PEVs (pressure equalization valves) remaining open and under electrical control;
  • reconfiguring DC1 & SM PkhO transfer tunnel and Russian segment (RS) systems; checking out BSS Orlan suit interface control panels in DC1 & PkhO;
  • reconfiguring EVA communications links;
  • rerouting C&W (caution & warning) alarms from the C&W panel to the PkhO EVA support panel;
  • checking out comm & medical monitoring gear as well as BK-3 repress bottles in the DC1;
  • tearing down the air duct in the DC1 (leaving the V3 fan place); and
  • conducting the final inspection of the Orlans and their BSS units (10:20am).

  After ISS attitude control was transferred to RS thrusters (for control during DC1 airlock depress), the crew received the Go for suit donning at 11:04am, initiating further events right on schedule.

Suit leak checks at 600 mmHg/Torr were nominal and continued to be OK with further depressurization.  At 11:49am, purging of the Orlans was completed, and prebreathe for denitrogenation began.  Egress occurred at 12:51pm, with both spacesuits working nominally.

After ingress from the EVA and DC1 airlock repressurization from cabin air at ~7:00pm, the crew will open hatches and reenter the SM for closeout operations.  Padalka first installs the DC1 air duct, then reactivates systems in the DC1 and other RS modules for restoration to pre-EVA conditions.

At ~7:30pm, Mike will begin reopening of the USOS transfer hatches from the RS.  Subsequently, the FE will restore the TCS and the OpsLAN.

A cabin air repress from Progress 15 is scheduled at ~8:00pm to replace the pressure drop from the DC1 repress.

At 10:25pm tonight, ISS attitude control will be handed over to RS MCS for maneuvering the station from the current (EVA-required) LVLH back to XPOP (x-axis perpendicular to orbit plane) at 10:35pm (yaw: -179.5 deg, pitch: -7 deg, roll: 0 deg).

Shifted sleep time begins late tonight at 12:00am/midnight, to extend till 9:00am tomorrow morning, to "ease" the crew back to their nominal sleep cycle.

More information and background material on EVA 4 can be found at the Expedition 9 EVA 4 Reference Page

Major upcoming events:

  • Reboost -- 9/22 (phase angle correction for 9S)
  • Soyuz 9S launch -- 10/9;
  • Soyuz 9S dock -- 10/11;
  • Soyuz 8S undock/land -- 10/19;
  • Soyuz 9S relocate -- 11/18;
  • Progress 16P launch -- 11/24.

Expedition 9 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

No CEO (crew earth observation) activities today.

Previous CEO images can be viewed at these websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:30pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is Off.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is Off.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is Off.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off. SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  •  
  • SM Working Compartment:  Pressure (mmHg) -- 762; temperature (deg C) -- 24.0; ppO2 (mmHg) -- 164.0; ppCO2 (mmHg) -- 5.0.
  • SM Transfer Compartment:  Pressure (mmHg) -- 752; temperature (deg C) -- 19.9.
  • FGB Cabin:  Pressure (mmHg) -- 756; temperature (deg C) -- 23.3.
  • Node:  Pressure (mmHg) -- 764.9; temperature (deg C) -- 24.8 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 767.1; temperature (deg C) -- 23.5; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) -- 766.9; temperature (deg C) -- 24.7; shell heater temp (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • (n/a = data not available)

  Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, "sun slicer", i.e., drag reduction-biased by 47 deg angle (2B: +47, 4B: -47).
  • SM batteries:  All batteries (8) are on line in "Partial Charge" mode.  
  • FGB batteries:  Battery #3 is off line.  All other batteries (5) are on line in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Discharge mode; PCU-2 is in Discharge mode.

  Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

  Propulsion System (PS):

  • Total propellant load available: 4220 kg (9303 lb) as of 9/2/04;  [SM(552) + FGB(3388) + Progress M(280)].  (Capabilities: SM -- 860 kg; FGB -- 6120 kg).

  Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

  Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 180.5 deg, pitch: -6.9 deg., roll: 0 deg]), with CMG TA (thruster assist) until 9/2 for EVA-11, then back to XPOP until next reboost (9/22).

  Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

  Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

 

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of this morning, 8:35am EDT [= epoch]):

  • Mean altitude -- 360.9 km
  • Apogee height -- 364.7 km
  • Perigee height -- 357.1 km
  • Period -- 91.76 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.000569
  • Solar Beta Angle -- 46.0 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.69
  • Mean altitude loss in last 24 hours -- 110 m
  • Revolutions since FGB/Zarya launch (Nov. '98)  -- 33045

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter